OpenStudy (anonymous):

Is the limit of ln x as x approaches to infiniti 1?

OpenStudy (anonymous):

\[\lim_{x \rightarrow \infty} \ln x / x = (1/x)/1 = 1/x\] \[1/\infty = 0\] I used L'Hopital's rule

OpenStudy (anonymous):

Oh I just meant ln x not lnx/x. Thanks though

OpenStudy (anonymous):

oh if u went ln x then it is unbounded and will be infinity

OpenStudy (anonymous):

Attached is a plot of Log[x], "Ln[x]", from x = 0 through x=10^100

OpenStudy (anonymous):

a plot of 0 to 10^100 does not always demonstrate end behavior, inthis case it does....but not always

OpenStudy (anonymous):

Just offered as a "it seems to be the case...".